TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x05x23x18x29

Powering Tests of Sector 7-8: activity of Wednesday 23/05.

Today the activity was concentrated on Q5, which was released by MPP after analysis of the quench of Monday morning.

The following tests have been executed on Q5 (MANUAL execution):
  • ramp to nominal (4310 A) with 5 A/s ramp rate: quenched occurred at 4210 A on B2 aperture; preliminary analysis confirmed it was a training quench;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a plateau of 10 min was mantained for possible voltage drift recording;
  • keeping B1 power converter at nominal current, B2 power converter was lowered to 55% of the nominal current (2370 A); with this unbalanced configuration, a quench was driven by QPS generating an artificial quench signal from one current lead detector; test was successful;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a FPA was provoked from power converter, with an apparently too high ramp rate which triggered the quench detector.

Proposal for the activity of tomorrow.

Q4 and D2 will be tested from CCC.

Works towards powering the main circuits in the arc and the 600 A correctors (before reaching 1.9 K):

  • plastic bags on 600 A leads (action S. Kockorowski);
  • IST of QPS for which 1.9 K is not required (action R. Denz);
Once 1.9 K are reached:
  • ElQA for the three 13 kA circuits and some selected 600 A correctors (list to be provided by HCC; action D. Bozzini);
  • IST of QPS at 1.9 K (action R. Denz);
  • OK from PIC, PC, EE and Current Leads.

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-23 - MirkoPojer
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login